OpenStudy (anonymous):

Angle C is an inscribed angle of circle P. Angle C measures negative x plus 8 degrees and arc AB measures 12 x degrees. Find x.

OpenStudy (jdoe0001):

same as before

OpenStudy (nikato):

So <C= -x+8 And mAB=12x http://www.mathwarehouse.com/geometry/circle/angle-tangent-and-chord.php So <C= 1/2 (mAB) Substitr your value into the equation -x+8=1/2(12x) Can you solve for x from here?

OpenStudy (jdoe0001):

\(\bf \textit{incribed angle}=\cfrac{\textit{intercepted arc}}{2}\implies 2\textit{incribed angle}=\textit{intercepted arc} \\ \quad \\ -x+8=\cfrac{12x}{2} \)